Two railroad freight cars with masses 100 Mg and 150 Mg approach with equal speeds of 0.310 m/s. They collide, the lighter car rebounding opposite its original direction at 0.270 m/s .

Find the velocity of the heavier car after the collision. Assume the original direction of the lighter car is positive.

What fraction of the original kinetic energy was lost in this inelastic collision?

Express your answers using two significant figures.
Answer #1

PART A

v1*m1+v2*m2=const.


100e-3*0.31-150e-3*0.31 = -0.27*100e-3+v*150e-3.

v=0.07666 m/s


PART B

K1=0.5m1v1^2+0.5*m2v2^2= 0.0120 J.

K2= 4.0782*10-3 J.

K1-K2/K1=66%